OpenStudy (anonymous):

Find the vertices and foci of the hyperbola with equation (give me a second to write it)

OpenStudy (anonymous):

\[\frac{( x+4)^2 }{ 9 }- \frac{ (y-5)^2 }{ 16 } = 1\]

OpenStudy (shinalcantara):

Hyperbola Standard Equation: \[\frac{ (x-h)^{2} }{ a ^{2} } - \frac{ (y-k)^{2} }{ b ^{2} } = 1\] where (h,k) is the vertex coordinates

OpenStudy (anonymous):

That I know

OpenStudy (shinalcantara):

Then you'll have the vertex coordinate Foci distance = 2a

OpenStudy (anonymous):

Vertex is (-4,5) So the foci are 18 away?

OpenStudy (anonymous):

Wait no they are 6 away

OpenStudy (anonymous):

Because a is squared

OpenStudy (shinalcantara):

yeah it's 6..

OpenStudy (anonymous):

Okay so now what would I do?

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

well, do we agree that when the x parts zero out, the equation makes no sense? 0 - y^2 = 1

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

so the verts are when y=5, since 5-5=0 the x parts are then solved for: (x+4)^2 = 9

OpenStudy (amistre64):

the foci of a hyperB are along the same line as the verts .... so they end in ,5 they also rely on the pythag thrm of the denominators

OpenStudy (anonymous):

Okay

OpenStudy (amistre64):

tell me what youhave so far

OpenStudy (anonymous):

That the vertices will end in 5

OpenStudy (amistre64):

verts and focuses, yes the x part for the verts must satisfy: (x+4)^2/9=1 so solve for x+4 = 3, and x+4 = -3

OpenStudy (amistre64):

the focuses we rely on the center, and a distance gotten from pythaging the denominators

OpenStudy (anonymous):

Oh so the vertices are (-1,5)(-7,5)

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

lets say |f|, the distance a focus is from the center, is equal to: |f| = sqrt(9+16) our center is -4 so our x parts for the focuses is: -4 +- sqrt(25)

OpenStudy (anonymous):

Got it Thanks so much, I really appreciate it

OpenStudy (amistre64):

youre welcome, and good luck

OpenStudy (anonymous):

Thank you :)